OpenStudy (anonymous):

find the limit L. then find delta>0 such that lf(x) - Ll < 0.01 whenever 0< lx-cl < delta. ----> lim (4 -(x/2)) x->4

OpenStudy (anonymous):

limit as \(x\to 4\) is \[4-2=2\] right? so you have \[|4-\frac{x}{2}-2|<0.01\]

OpenStudy (anonymous):

or \[|2-\frac{x}{2}|<0.01\] meaning \[|\frac{4-x}{2}|<0.01\]so \[|4-x|<0.02\] will work

OpenStudy (anonymous):

oh and it is clear that \(|4-x|=|x-4|\) right, so \(\delta=0.02\) will work for you here

OpenStudy (anonymous):

oh how do u prove that l4-xl=lx-4l ?

OpenStudy (anonymous):

in english, \(|x-4|\) is the distance between \(x\) and \(4\) so is \(|4-x|\)

OpenStudy (anonymous):

oh so can i show my work like this.... lx-4l=l -1(x-4) l

OpenStudy (anonymous):

you can check it if you like, but there is not much to prove. if \(a-b>0\) then \(|a-b|=a-b\) and \(|b-a|=-(b-a)=a-b\) as well you can see what happens if \(a-b<0\)

OpenStudy (anonymous):

i would take it as a given, and not bother proving it for this problem i would assume it

OpenStudy (anonymous):

how bout if i'm trying to solve l (-x/2) +2 l < 0.01

OpenStudy (anonymous):

can i do this: l (-1/2)(x-4) l < 0.01 and then (1/2) lx-4l < 0.01

OpenStudy (anonymous):

? :/

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

or \[|2-\frac{x}{2}|=|\frac{4-x}{2}|=\frac{1}{2}|4-x|\] any way you like

OpenStudy (anonymous):

oh yeah but that's if you're trying to match the absolute value to l4-xl....?